OpenStudy (anonymous):

f(x)=|2x-8|, find the value of f(1)-f(2)

OpenStudy (anonymous):

f(1) = I 2(1) - 8I = I 2 - 8I = I -6 I = 6 f(2) I 4 -8I = 4 6 - 4 = 2.